# Universal bound on the Lyapunov spectrum of quantum master equations

### v1

@@ 2.2 Quantum master equations
where now $\mathrm{Id}_{k}$ is the identity on $\mathcal{M}_{k}(\mathbb{C})$ and the inequality holds for all positive matrices $\tilde{\operatorname{O}}$ in $\mathcal{M}_{kd}(\mathbb{C})$. If we then denote by $\mathcal{P}_{k}$ the set of $k$-positive maps, it is rather straightforward to prove the chain of inclusions [ChrD2022]
$\mathcal{P}\,\equiv\,\mathcal{P}_{\mathfrak{1}}\supset\mathcal{P}_{\mathfrak{2}}\supset\dots\supset\mathcal{P}_{d}\,\equiv\,\mathcal{CP}$ (12)
@@
We thus see that the class of $d$-positive maps, i.e., of completely positive maps ($\mathcal{CP}$), is the smallest set in the above chain of inclusions. Unfortunately, for $k$-positive maps there is no general counterpart of the operator-sum representation (8) that characterizes completely positive maps. Analogously, on the differential level the conditions (7) have no general counterpart to specify generators of $k$-positive semigroups [WoCi2008, ChrD2022]. Even if it can be argued that completely positive maps are the only ones adapted to describe the evolution of open quantum systems (see, however, the debate [PecP1994, AliR1995, PecP1995] and the more recent [DoLi2016]), completely positive dynamical maps can be solutions of a quantum master equation (5) whose generator does not satisfy the complete positivity conditions (7). The upshot is that quantum master equations not satisfying (7) must be necessarily considered. We refer to [SmVa2010] for an example with clear physical interpretation. In addition, understanding the properties of $k$-positive semigroups for all $k$ is necessary to determine criteria to discriminate between entangled and separable states (see [ChKo2009] and references therein).
@@
We conclude the section recalling that a trace and self-adjointness preserving linear map $\Phi$ is positive if and only if is a contraction [KosA1972]
@@
$\left\|\Phi(\operatorname{O})\right\|_{\mathfrak{1}}\,\leq\,\left\|\operatorname{O}\right\|_{\mathfrak{1}}\qquad\forall\,\operatorname{O}=\operatorname{O}^{\dagger}\in\mathcal{M}_{d}(\mathbb{C})$
with respect the trace norm (also known as Schatten $1$-norm)
$\left\|\operatorname{O}\right\|_{\mathfrak{1}}:=\operatorname{Tr}\left(\sqrt{\operatorname{O}^{\dagger}\operatorname{O}}\right)\,.$
@@
The trace norm for self-adjoint operators coincides with the sum over the absolute value of all eigenvalues. Thus, positive linear maps are contractive and transform a convex compact set into itself. Again, these facts permit to demonstrate that positive maps satisfy the hypotheses of the Perron-Frobenius theorem. The conclusion is that all the eigenvalues belong to the unit disk of the complex plane and are symmetric with respect with the real axis. Furthermore, there is at least one eigenvalue equal to one whose right eigenoperator is a state operator under suitable normalization.
@@ 3 Embedded classical master equation
$\bm{\alpha}(0)=\sum_{i=1}^{d}p_{i}\,\bm{\varphi}_{i}\bm{\varphi}_{i}^{\dagger}$ (13)
@@
$\bm{\alpha}(t)=\sum_{i=1}^{d}\wp_{i}(t)\,\bm{\phi}_{i}(t)\bm{\phi}_{i}^{\dagger}(t)$ (14)
where both the $\bm{\varphi}_{i}$’s and the $\bm{\phi}_{\mathfrak{i}}(t)$’s constitute a complete orthonormal basis of $\mathbb{C}^{d}$ and the $p_{i}$’s and $\wp_{i}$’s respectively denote the eigenvalues of $\bm{\alpha}(0)$, $\bm{\alpha}(t)$.
@@
At any time $t$ the $d\times d$ matrices $\mathsf{W}(t)$ have the same structure as the generator of the classical master equation (2): they are now column-diagonally dominated with elements
$\mathsf{W}_{i,j}(t)=\left\langle\,\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\,,\mathfrak{L}(t)[\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)]\,\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\,\right\rangle=\begin{cases}\mathsf{R}_{i,j}(t)&\forall\,i\neq j\\[8.5359pt] -\sum_{k\neq i}^{d}\mathsf{R}_{i,j}(t)&\forall\,i=j\end{cases}$ (17)
@@
Before we continue some remarks are in order.
  • In general the “spectral flow” equations (15), (16) do not provide a constructive method to solve (5) but are only a tool to analyze the kinematic properties of individual solutions. More precisely, the matrix elements of (18) depend upon the unitary evolution operator in (15). The only general method to identify this latter one requires having first solved (5) for a given self-adjoint initial condition (13). This also means that the family of matrices $\mathsf{U}(t)$ and $\mathsf{W}(t)$ depend upon the choice of initial conditions. In other words, the spectral flow equations play for the kinematic analysis of quantum master equations the same role as the current velocity formalism for the analysis of the Fokker-Planck equation [NelE2001, § 13].
@@
    $\overline{\mathscr{x}(z)}=\mathscr{x}(\overline{z})\qquad\forall\,z\in\mathbb{C}\quad\text{ where }\operatorname{Re}z=t\,.$
@@
    $\bm{\alpha}^{\dagger}(z)=\bm{\alpha}(\bar{z})$ (19)
    in a neighborhood of the real $t$-axis. This fact allows us to invoke Theorem 6.1 in [KatT1982, § 2.6.1] guaranteeing that eigenvalues and spectral projectors of a family of operators enjoying (19) are also holomorphic functions and as such are continuous and differentiable. In addition [KatT1982, § 2.6.1], eigenvectors can also be chosen to be holomorphic functions, thus ensuring the validity of the construction used in the proof.
@@ 3.1 Kossakowski conditions
$\left\|\bm{\alpha}(t)\right\|_{\mathfrak{1}}=\sum_{i=1}^{d}|\wp_{i}(t)|=\left\|\bm{\wp}(t)\right\|_{\mathfrak{1}}$ (20)
@@
It is readily clear that if the complete positivity conditions (7) holds true then the off-diagonal entries (18) are positive functions of time thus implying that (16) is a classical master equation, and, as such describes contractive dynamics. Conversely, if the (18) are positive functions of time for any solution of (5) corresponding to self-adjoint initial data, then the semigroup generated by (5) is necessarily contractive and therefore positive [ChrD2022]. In fact, requiring that the (18) are always positive is equivalent to the Kossakowski conditions [KosA1972, GoKoSu1976] (see also [RiHu2012, § 4.3]) asserting that a linear map of the form (6) is the generator of a trace-preserving contraction semigroup if and only if its matrix entries evaluated on an arbitrary orthonormal frame specify a classical generator. More explicitly, if we denote by $\big\{\operatorname{P}_{i}\big\}_{i=1}^{d}$ a complete set of orthonormal projectors on $\mathbb{C}^{d}$ (that is,
$\sum_{i=1}^{d}\operatorname{P}_{i}=\operatorname{1}_{d}\&$ $\operatorname{P}_{i}\operatorname{P}_{j}=\delta_{i,j}\operatorname{P}_{i}\qquad\forall\,i,j=1,\dots,d\,)$
@@
$\begin{cases}\operatorname{Tr}\left(\operatorname{P}_{i}\mathfrak{L}(t)[\operatorname{P}_{j}]\right)\,\geq\,0&i\neq j\\[8.5359pt] \operatorname{Tr}\left(\operatorname{P}_{i}\mathfrak{L}(t)[\operatorname{P}_{i}]\right)=-\sum_{k\neq i=1}^{d}\operatorname{Tr}\left(\operatorname{P}_{k}\mathfrak{L}(t)[\operatorname{P}_{i}]\right)\,\leq\,0&i=j\end{cases}$
To summarize: the representation of solutions of the matrix equation (5) in terms of the eigenvector (15) and eigenvalue flow (15) equations provide us with implicit conditions to assess whether (5) generates a positive semigroup.
@@
We now turn to derive some properties of (16) that follow if we assume that the semigroup, in addition to being positive and trace-preserving, enjoys further properties. To this end, we recall that for any a complete orthonormal basis $\big\{\bm{f}_{i}\big\}_{i=1}^{d}$ of $\mathbb{C}^{d}$ the generator of a positive semigroup also satisfies
$\operatorname{Tr}\mathfrak{L}(t):=\sum_{i,j=1}^{d}\left\langle\,\bm{f}_{i}\,,\mathfrak{L}(t)[\bm{f}_{i}\bm{f}_{i}^{\dagger}]\,\bm{f}_{j}\,\right\rangle=\operatorname{Tr}\mathscr{L}(t)=-d\,\sum_{\ell=1}^{d^{2}-1}\mathscr{c}_{\ell}(t)\,\leq\,0$ (21)
which implies contractivity of the volume element for any $t\,\geq\,s$:
@@ 4.1 Condition on the embedded classical master equation imposed by $2$-positivity
$\bm{X}\colon\mathcal{M}_{d}(\mathbb{C})\mapsto\mathcal{M}_{d}(\mathbb{C})$
@@
$\operatorname{O}\,\geq\,0\quad\Longrightarrow\quad(\mathrm{Id}_{2}\otimes\bm{X})(\operatorname{O})\,\geq\,0\,.$
Let us now assume that (5) generates $2$-positive semigroup. Then we claim that the classical master equation constructed out of any of the particular solutions is such that
@@ 4.1.1 Proof of the claim
The action of a semigroup element close to the identity on a state operator $\bm{\rho}$ admits the explicit representation
$\Phi(t+\varepsilon,t)[\bm{\rho}]=\bm{\rho}+\varepsilon\,\mathfrak{L}(t)[\bm{\rho}]+O(\varepsilon^{2})$
If we pick a vector in $\mathbb{C}^{2d}$ of the form
$\bm{u}=\frac{\mathbf{e}_{1}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}^{\scriptscriptstyle{(i)}}(t)+\mathbf{e}_{2}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}^{\scriptscriptstyle{(j)}}(t)}{\sqrt{2}}$
where $\left\{\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\right\}_{i=1}^{d}$ is a complete orthonormal basis of $\mathbb{C}^{d}$ determined by the eigenvector flow equation (15). The projector along $\bm{u}$ is clearly a positive matrix
$\operatorname{P}_{u}=\bm{u}\bm{u}^{\dagger}=\frac{1}{2}\begin{bmatrix}\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)&\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)\\ \bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)&\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)\end{bmatrix}$.
@@
$(\mathrm{Id}_{2}\otimes\Phi(t+\varepsilon,t))[\operatorname{P}_{u}]$ $=\frac{1}{2}\begin{bmatrix}\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)&\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)\\ \bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)&\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)\end{bmatrix}$
$+\frac{\varepsilon}{2}\begin{bmatrix}\mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)\big]&\mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)\big]\\ \mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)\big]&\mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)\big]\end{bmatrix}+O(\varepsilon^{2})\,.$
@@
$\left\langle\,\bm{v}\,,(\mathrm{Id}_{2}\otimes\Phi(t+\varepsilon,t))[\operatorname{P}_{u}]\,\bm{v}\,\right\rangle\,\geq\,0$ $\forall\,\bm{v}\in\,\mathbb{C}^{2d}$
@@
$\bm{v}=\frac{\mathbf{e}_{1}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}^{\scriptscriptstyle{(i)}}(t)-\mathbf{e}_{2}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}^{\scriptscriptstyle{(j)}}(t)}{\sqrt{2}}=\frac{1}{\sqrt{2}}\begin{bmatrix}\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\\ -\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\end{bmatrix}$
@@
$\left\langle\,\bm{u}\,,\bm{v}\,\right\rangle=0\,.$
In such a case, after a straightforward calculation we obtain
$0\,\leq\,\lim_{\varepsilon\searrow 0}\frac{\left\langle\,\bm{v}\,,(\mathrm{Id}_{2}\otimes\Phi(t+\varepsilon,t))[\operatorname{P}_{u}]\bm{v}\,\right\rangle}{\varepsilon}$
$\hskip 14.22636pt=\frac{\left\langle\,\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)\big]\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\,\right\rangle+\left\langle\,\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)\big]\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\,\right\rangle}{2}$
$\hskip 14.22636pt-\frac{\left\langle\,\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\bm{\phi}^{\scriptscriptstyle{(j)}}{}^{\dagger}(t)\big]\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\,\right\rangle+\left\langle\,\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}^{\scriptscriptstyle{(j)}}(t)\bm{\phi}^{\scriptscriptstyle{(i)}}{}^{\dagger}(t)\big]\bm{\phi}^{\scriptscriptstyle{(i)}}(t)\,\right\rangle}{2}$

### v2

@@ 2.2 Quantum master equations
where now $\mathrm{Id}_{k}$ is the identity on $\mathcal{M}_{k}(\mathbb{C})$ and the inequality holds for all positive matrices $\tilde{\operatorname{O}}$ in $\mathcal{M}_{kd}(\mathbb{C})=\mathcal{M}_{k}(\mathbb{C})\otimes\mathcal{M}_{d}(\mathbb{C})$. If we then denote by $\mathcal{P}_{k}$ the set of $k$-positive maps, it is rather straightforward to prove the chain of inclusions [ChrD2022]
$\mathcal{P}\,\equiv\,\mathcal{P}_{1}\supset\mathcal{P}_{2}\supset\dots\supset\mathcal{P}_{d}\,\equiv\,\mathcal{CP}$ (23)
@@
We thus see that the class of $d$-positive maps, i.e., of completely positive maps ($\mathcal{CP}$), is the smallest set in the above chain of inclusions. Unfortunately, for $k$-positive maps, there is no general counterpart of the operator-sum representation (10) that characterizes completely positive maps. Analogously, at the differential level, the conditions (9) have no general counterpart to specify generators of $k$-positive semigroups [WoCi2008, ChrD2022]. Even if it can be argued that completely positive maps are the only ones adapted to describe the evolution of open quantum systems (see, however, the debate [PecP1994, AliR1995, PecP1995] and the more recent [DoLi2016]), completely positive dynamical maps can be solutions of a quantum master equation (7) whose generator does not satisfy the complete positivity conditions (9). The upshot is that quantum master equations that do not satisfy (9) must necessarily be considered. We refer to [SmVa2010] for an example with a clear physical interpretation. In addition, the study of $k$-positive maps for each $k$ is relevant to criteria for discriminating between entangled and separable states (see [ChKo2009] and the references therein). To summarize:
  1. 1.
    (8) equipped with regularity conditions on the time dependence is the generator of a semigroup of trace and self-adjointness preserving linear maps or, equivalently, of infinitely divisible trace and self-adjointness preserving linear maps;
  2. 2.
    we call the semigroup trace preserving and positive if all its elements are linear maps that send the convex set of state operators into itself;
  3. 3.
    the properties under tensor product composition with the identity map filter the set of positive linear maps into the flag of nested subsets (23). Unfortunately, in general the resulting classification does not lead to an explicit representation;
  4. 4.
    however, when (9) holds true the semigroup solution of a quantum master equation only consists of completely positive maps enjoying the property (21). This property is equivalent [StiW1955, ChoM1975] to the Kraus positive operator sum (10);
  5. 5.
    in the time-autonomous case the semigroup enjoys the time translational invariance property (14): it thus reduces to a one parameter family of linear maps.
@@
We conclude the section by recalling that a trace and self-adjointness preserving linear map $\Phi$ is positive if and only if it is a contraction [KosA1972]
@@
$\left\|\Phi(\operatorname{O})\right\|_{1}\,\leq\,\left\|\operatorname{O}\right\|_{1}\qquad\forall\,\operatorname{O}=\operatorname{O}^{\dagger}\in\mathcal{M}_{d}(\mathbb{C})$ (24)
with respect to the trace norm (also known as Schatten $1$-norm)
$\left\|\operatorname{O}\right\|_{1}:=\operatorname{Tr}\left(\sqrt{\operatorname{O}^{\dagger}\operatorname{O}}\right)$ (25)
@@
The trace norm for self-adjoint operators coincides with the sum of the absolute values of all eigenvalues. Thus, positive linear maps are contractive and transform a convex compact set into itself. Again, these facts allow us to demonstrate that positive maps satisfy the hypotheses of the Perron-Frobenius theorem. The conclusion is that all the eigenvalues belong to the unit disk of the complex plane and are symmetric with respect to the real axis. Furthermore, there is at least one eigenvalue equal to one. The right eigenoperator associated with this eigenvalue is a state operator after appropriate normalization.
@@ 3 Embedded classical master equation
$\bm{\rho}(0)=\sum_{i=1}^{d}p_{i}\,\bm{\varphi}_{i}\bm{\varphi}_{i}^{\dagger}$ (26)
@@
$\bm{\rho}(t)=\sum_{i=1}^{d}\wp_{i}(t)\,\bm{\phi}_{i}(t)\bm{\phi}_{i}^{\dagger}(t)$ (27)
where both the $\bm{\varphi}_{i}$’s and the $\bm{\phi}_{i}(t)$’s constitute a complete orthonormal basis of $\mathbb{C}^{d}$ and the $p_{i}$’s and $\wp_{i}$’s respectively denote the eigenvalues of $\bm{\rho}(0)$, $\bm{\rho}(t)$.
@@
At any time $t$ the $d\times d$ matrices $\mathsf{W}(t)$ have the same structure as the generator of the classical master equation (2): now they are column-diagonally dominated with elements
$\mathsf{W}_{i,j}(t)=\left\langle\,\bm{\phi}_{i}(t)\,,\mathfrak{L}(t)[\bm{\phi}_{j}(t)\bm{\phi}_{j}^{\dagger}(t)]\,\bm{\phi}_{i}(t)\,\right\rangle=\begin{cases}\mathsf{R}_{i,j}(t)&\forall\,i\neq j\\[8.5359pt] -\sum_{k\neq j}^{d}\mathsf{R}_{k,j}(t)&\forall\,i=j\end{cases}$ (30)
@@
Before continuing, some remarks are in order.
  • In general, the “spectral flow” equations (28), (29) do not provide a constructive method for solving (7) but are only a tool to analyze the kinematic properties of individual solutions. More precisely, the matrix elements of (31) depend on the unitary evolution operator in (28). The only general method to identify the latter requires first solving (7) for a given self-adjoint initial condition (26). This also means that the families of matrices $\mathsf{U}(t)$ and $\mathsf{W}(t)$ depend on the choice of the initial conditions. In other words, the spectral flow equations play the same role for the kinematic analysis of quantum master equations as the current velocity formalism plays for the kinematic analysis of the Fokker-Planck equation [NelE2001, § 13].
@@
    $\overline{\mathscr{x}(z)}=\mathscr{x}(\overline{z})\qquad\forall\,z\in\mathbb{C}\quad\text{ where }\operatorname{Re}z=t\,.$ (32)
@@
    $\bm{\rho}^{\dagger}(z)=\bm{\rho}(\bar{z})$ (33)
    in a neighborhood of the real $t$-axis. This fact allows us to invoke Theorem 6.1 in [KatT1982, § 2.6.1], which guarantees that the eigenvalues and spectral projectors of a family of operators satisfying (33) are holomorphic functions and, as such, are continuous and differentiable. In addition [KatT1982, § 2.6.1], eigenvectors can also be chosen to be holomorphic functions, thus ensuring the validity of the construction used in the proof.
@@ 3.1 Kossakowski conditions
$\left\|\bm{\rho}(t)\right\|_{1}=\sum_{i=1}^{d}|\wp_{i}(t)|=\left\|\bm{\wp}(t)\right\|_{1}$ (34)
@@
It is readily clear that if the complete positivity conditions (9) hold true, then the off-diagonal entries (31) are positive functions of time. Consequently, (29) is a classical master equation and, as such, describes contractive dynamics. Conversely, if the entries (31) are always positive for any solution of (7) corresponding to self-adjoint initial data, then the semigroup generated by (7) is necessarily contractive and, therefore, positive [ChrD2022]. In fact, requiring that the (31) are always positive is equivalent to the Kossakowski conditions [KosA1972, GoKoSu1976] (see also [RiHu2012, § 4.3]) asserting that a linear map of the form (8) is the generator of a trace-preserving contraction semigroup if and only if its matrix entries evaluated on an arbitrary orthonormal frame specify a classical generator. More explicitly, if we denote by $\big\{\operatorname{P}_{i}\big\}_{i=1}^{d}$ a complete set of orthonormal projectors on $\mathbb{C}^{d}$, that is,
$\begin{split}&\sum_{i=1}^{d}\operatorname{P}_{i}=\operatorname{1}_{d}\\ &\operatorname{P}_{i}\operatorname{P}_{j}=\delta_{i,j}\operatorname{P}_{i}\end{split}$ $\qquad\forall\,i,j=1,\dots,d$ (35)
@@
$\begin{cases}\operatorname{Tr}\left(\operatorname{P}_{i}\mathfrak{L}(t)[\operatorname{P}_{j}]\right)\,\geq\,0&i\neq j\\[8.5359pt] \operatorname{Tr}\left(\operatorname{P}_{i}\mathfrak{L}(t)[\operatorname{P}_{i}]\right)=-\sum_{k\neq i=1}^{d}\operatorname{Tr}\left(\operatorname{P}_{k}\mathfrak{L}(t)[\operatorname{P}_{i}]\right)\,\leq\,0&i=j\end{cases}$ (36)
To summarize: the identity (34) allows us to explicitly relate the contractivity of the semigroup generated by (7) to that of solutions of the classical master equation (29) and, therefore, to derive Kossakowski’s conditions.
@@
We now turn to derive some properties of (29) that follow if we assume that the semigroup, in addition to being trace preserving and positive, enjoys further properties. To this end, we recall that the Hilbert-Schmidt inner product of two matrices is defined by
$\langle\operatorname{A},\operatorname{B}\rangle:=\operatorname{Tr}\!\left(\operatorname{A}^{\dagger}\operatorname{B}\right)\qquad\forall\,\operatorname{A},\operatorname{B}\in\mathcal{M}_{d}(\mathbb{C})$ (37)
Accordingly, the trace of $\mathfrak{L}(t)$ is
$\operatorname{Tr}\mathfrak{L}(t):=\sum_{\mu=1}^{d^{2}}\langle\operatorname{A}_{\mu}\mathfrak{L}(t)[\operatorname{A}_{\mu}]\rangle$ (38)
where $\{\operatorname{A}_{\mu}\}_{\mu=1}^{d^{2}}$ is a complete orthonormal basis of $\mathcal{M}_{d}(\mathbb{C})$. In particular, for any complete orthonormal basis $\{\bm{f}_{i}\}_{i=1}^{d}$ of $\mathbb{C}^{d}$, the collection of outer products $\{\bm{f}_{i}\bm{f}_{j}^{\dagger}\}_{i,j=1}^{d}$ constitutes a complete orthonormal basis of $\mathcal{M}_{d}(\mathbb{C})$. Hence, we obtain
$\operatorname{Tr}\mathfrak{L}(t)$ $=\sum_{i,j=1}^{d}\operatorname{Tr}\!\left((\bm{f}_{i}\bm{f}_{j}^{\dagger})^{\dagger}\mathfrak{L}(t)[\bm{f}_{i}\bm{f}_{j}^{\dagger}]\right)$
$=\sum_{i,j=1}^{d}\left\langle\bm{f}_{i},\mathfrak{L}(t)[\bm{f}_{i}\bm{f}_{j}^{\dagger}]\,\bm{f}_{j}\right\rangle$ (39)
It is straightforward to verify that (39) is also equal to the trace of the reshaped generator $\mathscr{L}(t)$ defined by (19):
$\operatorname{Tr}\mathfrak{L}(t)=\operatorname{Tr}\mathscr{L}(t)=-d\,\sum_{\ell=1}^{d^{2}-1}\mathscr{c}_{\ell}(t)\,\leq\,0$ (40)
This result implies contractivity of the volume element for any $t\,\geq\,s$:
@@ 4.1 Condition on the embedded classical master equation imposed by 2-positivity
$\bm{X}\colon\mathcal{M}_{d}(\mathbb{C})\mapsto\mathcal{M}_{d}(\mathbb{C})$ (42)
@@
$\operatorname{O}\,\geq\,0\quad\Longrightarrow\quad(\mathrm{Id}_{2}\otimes\bm{X})(\operatorname{O})\,\geq\,0\,.$ (43)
Now, let us assume that (7) generates a $2$-positive semigroup. Then we claim that the classical master equation constructed out of any of the particular solutions is such that
@@ 4.1.1 Proof of the claim
The action of a semigroup element close to the identity on a state operator $\bm{\rho}$ admits the explicit representation:
$\Phi(t+\varepsilon,t)[\bm{\rho}]=\bm{\rho}+\varepsilon\,\mathfrak{L}(t)[\bm{\rho}]+O(\varepsilon^{2})$ (45)
Let us denote by $\mathbf{e}_{i}^{\scriptscriptstyle{(2)}}$, $i=1,2$ the elements of the canonical (or computational) basis of $\mathbb{C}^{2}$. We can then construct a vector in $\mathbb{C}^{2d}$ of the form
$\bm{u}=\frac{\mathbf{e}_{1}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}_{i}(t)+\mathbf{e}_{2}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}_{j}(t)}{\sqrt{2}}$ (46)
where $\left\{\bm{\phi}_{i}(t)\right\}_{i=1}^{d}$ is a complete orthonormal basis of $\mathbb{C}^{d}$ determined by the eigenvector flow equation (28). The projector along $\bm{u}$ is clearly a positive matrix
$\operatorname{P}_{u}=\bm{u}\bm{u}^{\dagger}=\frac{1}{2}\begin{bmatrix}\bm{\phi}_{i}(t)\bm{\phi}_{i}^{\dagger}(t)&\bm{\phi}_{i}(t)\bm{\phi}_{j}^{\dagger}(t)\\ \bm{\phi}_{j}(t)\bm{\phi}_{i}^{\dagger}(t)&\bm{\phi}_{j}(t)\bm{\phi}_{j}^{\dagger}(t)\end{bmatrix}$. (47)
@@
$(\mathrm{Id}_{2}\otimes\Phi(t+\varepsilon,t))[\operatorname{P}_{u}]$ $=\frac{1}{2}\begin{bmatrix}\bm{\phi}_{i}(t)\bm{\phi}_{i}^{\dagger}(t)&\bm{\phi}_{i}(t)\bm{\phi}_{j}^{\dagger}(t)\\ \bm{\phi}_{j}(t)\bm{\phi}_{i}^{\dagger}(t)&\bm{\phi}_{j}(t)\bm{\phi}_{j}^{\dagger}(t)\end{bmatrix}$
$+\frac{\varepsilon}{2}\begin{bmatrix}\mathfrak{L}(t)\big[\bm{\phi}_{i}(t)\bm{\phi}_{i}^{\dagger}(t)\big]&\mathfrak{L}(t)\big[\bm{\phi}_{i}(t)\bm{\phi}_{j}^{\dagger}(t)\big]\\ \mathfrak{L}(t)\big[\bm{\phi}_{j}(t)\bm{\phi}_{i}^{\dagger}(t)\big]&\mathfrak{L}(t)\big[\bm{\phi}_{j}(t)\bm{\phi}_{j}^{\dagger}(t)\big]\end{bmatrix}+O(\varepsilon^{2})\,.$ (48)
@@
$\left\langle\,\bm{v}\,,(\mathrm{Id}_{2}\otimes\Phi(t+\varepsilon,t))[\operatorname{P}_{u}]\,\bm{v}\,\right\rangle\,\geq\,0$ $\forall\,\bm{v}\in\,\mathbb{C}^{2d}$ (49)
@@
$\bm{v}=\frac{\mathbf{e}_{1}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}_{i}(t)-\mathbf{e}_{2}^{\scriptscriptstyle{(2)}}\otimes\bm{\phi}_{j}(t)}{\sqrt{2}}=\frac{1}{\sqrt{2}}\begin{bmatrix}\bm{\phi}_{i}(t)\\ -\bm{\phi}_{j}(t)\end{bmatrix}$ (50)
@@
$\left\langle\,\bm{u}\,,\bm{v}\,\right\rangle=0\,.$ (51)
In such a case, after a straightforward calculation, we obtain
$0\,\leq\,\lim_{\varepsilon\downarrow 0}\frac{\left\langle\,\bm{v}\,,(\mathrm{Id}_{2}\otimes\Phi(t+\varepsilon,t))[\operatorname{P}_{u}]\bm{v}\,\right\rangle}{\varepsilon}$
$\hskip 14.22636pt=\frac{\left\langle\,\bm{\phi}_{i}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}_{i}(t)\bm{\phi}_{i}^{\dagger}(t)\big]\bm{\phi}_{i}(t)\,\right\rangle+\left\langle\,\bm{\phi}_{j}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}_{i}(t)\bm{\phi}_{i}^{\dagger}(t)\big]\bm{\phi}_{j}(t)\,\right\rangle}{2}$
$\hskip 14.22636pt-\frac{\left\langle\,\bm{\phi}_{i}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}_{i}(t)\bm{\phi}_{j}^{\dagger}(t)\big]\bm{\phi}_{j}(t)\,\right\rangle+\left\langle\,\bm{\phi}_{j}(t)\,,\mathfrak{L}(t)\big[\bm{\phi}_{j}(t)\bm{\phi}_{i}^{\dagger}(t)\big]\bm{\phi}_{i}(t)\,\right\rangle}{2}$ (52)